# Stochastic Re-weighted Gradient Descent via Distributionally Robust Optimization

### v4

@@ Abstract
We present Re-weighted Gradient Descent (RGD), a novel optimization technique that improves the performance of deep neural networks through dynamic sample importance weighting. Our method is grounded in the principles of distributionally robust optimization (DRO) with Kullback-Leibler divergence. RGD is simple to implement, computationally efficient, and compatible with widely used optimizers such as SGD and Adam. We demonstrate the broad applicability and impact of RGD by achieving state-of-the-art results on diverse benchmarks, including improvements of +0.7% (DomainBed), +1.44% (tabular classification), +1.94% (GLUE with BERT), and +1.01% (ImageNet-1K with ViT).
Machine Learning, ICML
@@ 1 Introduction
Deep neural networks (DNNs) have become essential for solving a wide range of tasks, including image classification, object detection, machine translation, and speech recognition. The most commonly used paradigm for learning DNNs is empirical risk minimization (ERM Vapnik (1999)), which aims to identify a network that minimizes the average loss of training data points. Several algorithms, including SGD (Nemirovsky et al., 1983), Adam (Kingma & Ba, 2014), and Adagrad (Duchi et al., 2011), have been proposed for solving ERM. However, a drawback of ERM is that it weighs all the samples equally, often ignoring the rare and more difficult samples and focusing on the easier and abundant samples. This leads to suboptimal performance on unseen data, especially when the training data is scarce (Namkoong & Duchi, 2017). Consequently, recent works have developed data re-weighting techniques for improving the performance of ERM. One particularly fruitful approach in this line of work is the framework of Distributionally Robust Optimization (DRO) (Ben-Tal et al., 2013), which assigns higher weights to hard examples, often leading to models with better performance than ERM.
@@
$\sup_{P^{\prime}:D(P^{\prime}||P_{\text{data}})\leq\rho}\mathbb{E}_{P^{\prime}}[\ell(z,\theta)].$
Here, $\rho$ is the perturbation radius, and $D$ is a divergence that measures the distance between two probability distributions. Popular choices for $D$ include $f$-divergences, which are defined as $D_{f}(Q||P)=\mathbb{E}_{P}[f(dQ/dP)]$ for some convex function $f:\mathbb{R}_{+}\to\mathbb{R}$. In this case, Shapiro (2017) derived the following dual formulation of the above objective
@@
Here, $f^{*}(s)=\sup_{t}\{st-f(t)\}$ is the Fenchel conjugate of $f$. This alternative way of expressing DRO shows how it implicitly reweighs samples using the conjugate $f^{*}$. The seminal works of Duchi et al. (2016); Namkoong & Duchi (2017) studied this objective for various $f$-divergences and showed that it has variance reduction properties, and leads to models with good generalization performance under small perturbations ($\rho=O(1/n)$, where $n$ is the data size). Duchi & Namkoong (2018) showed that DRO under large perturbations ($\rho=O(1)$) leads to models with good fairness, tail risk guarantees. Li et al. (2020) (and its extended version Li et al. (2023)) considered KL divergence DRO, which is obtained by choosing $f(x)=x\log{x}$, and showed that setting dual variable $\lambda$ to a negative value results in robust models that can withstand corruptions in the training data.
Inspired by these properties, several recent studies have developed algorithms for optimizing DRO and designed data re-weighting techniques for various learning tasks. These algorithms fall into two broad categories: (a) Primal-Dual techniques which rely on alternating mirror ascent, descent to solve the min-max DRO objective (Namkoong & Duchi, 2016; Yan et al., 2020b; Fidon et al., 2020), and (b) Compositional optimization techniques which solve an equivalent compositional form of DRO, which takes the form $g(\mathbb{E}_{z}[h(z,\theta)])$, for some functions $g,h$ (Qi et al., 2021, 2020, 2022; Li et al., 2023). While these algorithms come with good convergence guarantees, they have certain drawbacks that limit their use in practice. (a) Scalability: primal-dual algorithms require updating and sampling from a probability distribution over the entire dataset at each iteration, making them computationally expensive. Although compositional optimization techniques alleviate this issue, gradient estimation within these algorithms is non-trivial and often necessitates maintaining moving averages of sample weights (Qi et al., 2020; Li et al., 2020). This introduces additional hyperparameters, complicating their application to large-scale scenarios. (b) Robustness to Outliers: many real-world datasets contain outliers, which pose challenges to algorithms optimizing DRO. In particular, these outliers often result in poor performance and instability during the DRO training process (Zhu et al., 2022; Zhai et al., 2021). Existing works often fail to account for outliers in real datasets, leading to subpar performance (see Table 1 for a detailed comparison).
@@
In this work, we address the aforementioned limitations of DRO optimization techniques. We focus on KL divergence-based DRO, and develop a lightweight algorithm for efficiently solving the resulting objective. Our algorithm simply optimizes the inner objective in Equation 1 using SGD. This gives rise to our stochastic Re-weighted Gradient Descent (RGD) algorithm, a variant of the classical SGD, that re-weights data points during each optimization step based on their difficulty. A key component of our algorithm is weight clipping that we introduce to protect against (benign) outliers and stabilize the algorithm. As demonstrated in our experiments (see Section 4, Appendix F), weight clipping significantly improves the performance of our algorithm on numerous learning tasks involving real-world datasets. Another noteworthy aspect of our algorithm is that it has the same runtime as SGD, has only one hyper-parameter, and scales to models with billions of parameters.
@@ 1.1 Evaluation
Supervised Learning: We evaluate RGD on several supervised learning tasks in language and vision domains. In the language domain, we apply RGD for BERT fine-tuning on the General Language Understanding Evaluation (GLUE) benchmark and show that RGD outperforms the BERT baseline by $+1.94\%$. In the vision domain, we apply RGD for ImageNet-1K classification using ViT-S model, and show that RGD outperforms the ViT-S baseline by $+1.01\%$.
Tabular Classification: Recently, Majmundar et al. (2022) introduced a tabular representation learning method called MET. Deep learning methods trained with the learned representations from MET achieved SOTA performance on downstream classification tasks, significantly improving upon Gradient Boosting decision trees (GBDT; Friedman (2001)). Our experiments show that applying RGD to the MET framework improves its performance by $1.51\%$ and $1.27\%$ on binary and multi-class tabular classification, respectively.
Domain Generalization: In domain generalization, the distributions of train and test datasets could be different (for example, training on pictures of real dogs and evaluating cartoon dogs). This task requires robustness to distribution shifts and DRO is a natural framework in this context. Gulrajani & Lopez-Paz (2020) showed that the ERM framework applied over deep networks, is highly effective for this problem. Perhaps surprisingly, this remained the state-of-the-art (SOTA) algorithm for a long time. Only recently, ERM has been beaten on this challenging task (Cha et al., 2022; Addepalli et al., 2022). In this work, we show that using RGD on top of these recent techniques further boosts their performance by $0.7\%$ and gives SOTA results on this task.
Meta-Learning: In meta-learning, we aim to learn models that generalize to new tasks with limited data. Predominant approaches in this domain use the classical ERM to optimize deep networks (Finn et al., 2017; Snell et al., 2017; Kumar et al., 2022). However, a common issue in this domain is that the learned models solve most tasks but fail catastrophically in some tasks. Consequently, this has promoted works that focus on worst-case performance (Collins et al., 2020). Recently, (Li et al., 2023) used KL-DRO to tackle this problem. However, the authors applied their algorithm for solving meta-regression on a toy-dataset that is free of outliers, and haven’t showcased their algorithm on practically relevant datasets such as Omniglot, miniImageNet. In this work, we show that using RGD as an off-the-hat addition to Model-Agnostic Meta-Learning (MAML) (Finn et al., 2017) can significantly improve the worst-case accuracy of these models on meta-classification benchmarks such as Omniglot, miniImageNet by up to $3\%$.
@@ 1.2 Contributions
  • State-of-the-Art (SOTA) Performance Enhancement. Extensive experiments demonstrate that RGD delivers significant performance gains across diverse learning tasks. In tabular classification, RGD boosts the accuracy of MET (Majmundar et al., 2022) by +1.44%. For out-of-domain generalization, RGD outperforms FRR (Addepalli et al., 2022) on DomainBed by +0.7%. Additionally, RGD improves the performance of BERT on GLUE benchmarks by +1.94% and ViT on ImageNet-1K by +1.01%. (see Section 4, Appendix F)
@@ 2 Related Work
DRO dates back to the early works of Ben-Tal et al. (2009, 2013). Since then several works have studied various statistical and optimization aspects of DRO. The seminal works of Lam (2016); Namkoong & Duchi (2017); Duchi et al. (2021) formally showed that minimizing empirical DRO risk - under small perturbations ($\rho=O(1/n)$, where $n$ is the data size) - is equivalent to minimizing sum of empirical risk and its standard deviation. Consequently, optimizing DRO leads to a better bias-variance trade-offs and generalizing models. In this work, we rely on this property of DRO to develop our re-weighting scheme. In another seminal work, Duchi & Namkoong (2018) showed that DRO - under large perturbations ($\rho=O(1)$) - leads to models with good tail performance.
@@
The aforementioned properties of DRO has led to numerous works applying it in various learning scenarios. For instance, Duchi & Namkoong (2018); Sagawa et al. (2019); Qi et al. (2021, 2020); Li et al. (2020, 2023) used DRO to tackle problems of class-imbalanaced classification and fairness. In another line of work, Namkoong & Duchi (2017); Fidon et al. (2020) studied DRO for designing models that generalize better than ERM. Our work falls in this second category of works that focus on generalization.
@@
From an optimization perspective, several works have focused on designing efficient algorithms for optimizing the DRO objective. These algorithms can be classified into two broad categories: primal-dual (Namkoong & Duchi, 2016; Fidon et al., 2020; Yan et al., 2020b), compositional optimization techniques (Qi et al., 2021, 2020, 2022; Li et al., 2020). One of the key drawbacks of primal-dual techniques is that they update and sample from a probability distribution over the entire training data at each step (aka. dual variables). A naive implementation of this step takes $O(n)$ time, which is prohibitive for large-scale tasks. Namkoong & Duchi (2016) reduced the complexity of this step to $O(\log{n})$ using data structures such as balanced binary search trees. However, the resulting algorithms are hard to implement in practice. Another drawback of these algorithms is that they require storing a buffer of weights for the entire dataset, which is infeasible at the scale of LLMs. Even if storing the weights is feasible, the presence of data augmentations complicates the resulting algorithms. Compositional optimization algorithms overcome these drawbacks by working with an equivalent formulation of DRO that can be written as composition of two functions: $g(\mathbb{E}_{z}[h(z,\theta)])$. One major drawback of these techniques though is that estimating the gradient $\nabla_{\theta}g(\mathbb{E}_{z}[h(z,\theta)])$ from a mini-batch is non-trivial. This requires certain additional steps in the algorithm which add to its complexity. For instance, the algorithm of Qi et al. (2021) requires making two backward passes at each step of SGD. The convergence analysis of Li et al. (2020) required two independent mini-batches at each iteration. The algorithms of Qi et al. (2020); Li et al. (2020) both require maintaining a moving average of the weights of the mini-batches, thus adding an additional hyper-parameter to the algorithm.
@@
Outlier robust DRO. Recent works have shown that DRO is highly sensitive to outliers (Zhai et al., 2021; Zhu et al., 2022). This is because DRO tends to magnify the influence of outliers by upweighting them further. This is one of the reasons for the suboptimal performance of existing DRO algorithms on real world datasets. To address this, Zhai et al. (2021) consider an adversarial model for outliers (where $\epsilon$-fraction of training data could be arbitrarily corrupted by a malicious adversary). They develop a heuristic to remove the outliers during each descent step for $\chi^{2}$-DRO and CVaR. While interesting, this is too strong of an adversary model which leads to data wastage in practice.
@@
Min-min DRO. Min-min DRO minimizes the following objective: $\inf_{P^{\prime}:D(P^{\prime}||P_{\text{data}})\leq\rho}\mathbb{E}_{P^{\prime}}[\ell(z,\theta)]$. Contrast this with DRO which minimizes: $\sup_{P^{\prime}:D(P^{\prime}||P_{\text{data}})\leq\rho}\mathbb{E}_{P^{\prime}}[\ell(z,\theta)]$. Unlike DRO which is primarily studied for generalization and fairness properties, min-min DRO is studied for training in the presence of outliers (Li et al., 2020; Kumar & Amid, 2021; Majidi et al., 2021). Instead of upweighting high loss points, min-min DRO downweights them.
@@
Other applications of DRO. Sagawa et al. (2019) optimized Group DRO for fair models when the group information is known. (Sinha et al., 2017) studied DRO with Wasserstein divergence for learning models that are robust to adversarial perturbations. DRO also appears in many classical statistical problems. For example, many boosting algorithms (including AdaBoost) can be viewed as performing DRO with KL-divergence-based uncertainty sets (Arora et al., 2012; Friedman, 2001). Faury et al. (2020) relied on KL-DRO for counterfactual risk minimization. Sakhi et al. (2020) use DRO for improving offline contextual bandits algorithms.

### v5

@@ Abstract
We present Re-weighted Gradient Descent (RGD), a novel optimization technique that improves the performance of deep neural networks through dynamic sample re-weighting. Leveraging insights from distributionally robust optimization (DRO) with Kullback-Leibler divergence, our method dynamically assigns importance weights to training data during each optimization step. RGD is simple to implement, computationally efficient, and compatible with widely used optimizers such as SGD and Adam. We demonstrate the effectiveness of RGD on various learning tasks, including supervised learning, meta-learning, and out-of-domain generalization. Notably, RGD achieves state-of-the-art results on diverse benchmarks, with improvements of +0.7% on DomainBed, +1.44% on tabular classification, +1.94% on GLUE with BERT, and +1.01% on ImageNet-1K with ViT.
@@ 1 Introduction
Deep neural networks (DNNs) have become essential for solving a wide range of tasks, including image classification, object detection, machine translation, and speech recognition. The most commonly used paradigm for learning DNNs is empirical risk minimization (ERM Vapnik (1999)), which aims to identify a network that minimizes the average loss of training data points. Several algorithms, including SGD (Nemirovsky et al., 1983), Adam (Kingma & Ba, 2015), and Adagrad (Duchi et al., 2011), have been proposed for solving ERM. However, a drawback of ERM is that it weighs all the samples equally, often ignoring the rare and more difficult samples and focusing on the easier and abundant samples. This leads to suboptimal performance on unseen data, especially when the training data is scarce (Namkoong & Duchi, 2017). Consequently, recent works have developed data re-weighting techniques for improving the performance of ERM. One particularly fruitful approach in this line of work is the framework of Distributionally Robust Optimization (DRO) (Ben-Tal et al., 2013), which assigns higher weights to hard examples, often leading to models with better performance than ERM.
@@
$\inf_{\theta}\sup_{P^{\prime}:D(P^{\prime}||P_{\text{data}})\leq\rho}\mathbb{E}_{P^{\prime}}[\ell(z,\theta)].$
Here, $\rho$ is the perturbation radius, and $D$ is a divergence that measures the distance between two probability distributions. Popular choices for $D$ include $f$-divergences, which are defined as $D_{f}(Q||P)=\mathbb{E}_{P}[f(dQ/dP)]$ for some convex function $f:\mathbb{R}_{+}\to\mathbb{R}$. In this case, Shapiro (2017) derived the following equivalent dual formulation of the above objective
@@
Here, $f^{*}(s)=\sup_{t}\{st-f(t)\}$ is the Fenchel conjugate of $f$. This alternative way of expressing DRO shows how it implicitly reweighs samples using the conjugate $f^{*}$. The seminal works of Duchi et al. (2021a); Namkoong & Duchi (2017) studied this objective for various $f$-divergences and showed that it has variance reduction properties, and leads to models with good generalization performance under small perturbations ($\rho=O(1/n)$, where $n$ is the data size). Furthermore, Duchi & Namkoong (2018) showed that DRO under large perturbations ($\rho=O(1)$) leads to models with good fairness, tail risk guarantees. In another line of work, Li et al. (2021) (and its extended version Li et al. (2023)) considered KL divergence DRO, which is obtained by choosing $f(x)=x\log{x}$, and showed that setting dual variable $\lambda$ to a negative value results in robust models that can withstand corruptions in the training data.
Inspired by these impressive properties, several recent studies have developed algorithms for optimizing DRO and designed data re-weighting techniques for various learning tasks. These algorithms fall into two broad categories: (a) Primal-Dual techniques which rely on alternating mirror ascent, descent to solve the min-max DRO objective (Namkoong & Duchi, 2016; Yan et al., 2020b; Fidon et al., 2020), and (b) Compositional optimization techniques which solve an equivalent compositional/dual form of DRO, which takes the form $g(\mathbb{E}_{z}[h(z,\theta)])$, for some functions $g,h$ (Qi et al., 2021; 2023b; 2023a; Li et al., 2023). While these algorithms come with good convergence guarantees, they have certain drawbacks that limit their use in practice. (a) Scalability: primal-dual algorithms require updating and sampling from a probability distribution over the entire dataset at each iteration, making them computationally expensive. Although compositional optimization techniques alleviate this issue, gradient estimation within these algorithms is non-trivial as the objective is no longer an empirical mean of the losses evaluated at the training data points. Overcoming this often necessitates maintaining moving averages of sample weights, which introduces additional hyperparameters, complicating their application to large-scale scenarios (Qi et al., 2023b; Li et al., 2021). (b) Robustness to Outliers: many real-world datasets contain outliers [^1] An outlier is a data point that lies significantly outside the typical pattern of a dataset. These outliers could be because of noise in data collection process or could be introduced by a malicious adversary. In this work, we are primarily concerned about the former type of outliers., which pose challenges to algorithms optimizing DRO. In particular, these outliers often result in poor performance and instability during the DRO training process (Zhu et al., 2022a; Zhai et al., 2021). Existing works often fail to account for outliers in real datasets, leading to subpar performance (see Table 1 for a detailed comparison).
@@
In this work, we address the aforementioned limitations of DRO optimization techniques. We focus on KL divergence-based DRO, and develop a lightweight algorithm for efficiently solving the resulting objective. Our algorithm simply optimizes the inner objective in Equation 1 using SGD. This gives rise to our stochastic Re-weighted Gradient Descent (RGD) algorithm, a variant of the classical SGD, that re-weights data points during each optimization step based on their difficulty. A key component of our algorithm is weight clipping that we introduce to protect against (benign) outliers and stabilize the algorithm. As demonstrated in our experiments (see Section 4, Appendix D), weight clipping significantly improves the performance of our algorithm on numerous learning tasks involving real-world datasets. Another noteworthy aspect of our algorithm is that it has the same runtime as SGD, has only one hyper-parameter, and scales to models with billions of parameters.
@@ 1.1 Evaluation
Supervised Learning:
We evaluate RGD on several supervised learning tasks in language and vision domains. In the language domain, we apply RGD for BERT fine-tuning on the General Language Understanding Evaluation (GLUE) benchmark and show that RGD outperforms the BERT baseline by $+1.94\%$. In the vision domain, we apply RGD for ImageNet-1K classification using ViT-S model, and show that RGD outperforms the ViT-S baseline by $+1.01\%$.
Tabular Classification:
Recently, Majmundar et al. (2022) introduced a tabular representation learning method called MET. Deep learning methods trained with the learned representations from MET achieved SOTA performance on downstream classification tasks, significantly improving upon Gradient Boosting decision trees (GBDT; Friedman (2001)). Our experiments show that applying RGD to the MET framework improves its performance by $1.51\%$ and $1.27\%$ on binary and multi-class tabular classification, respectively.
Domain Generalization:
In domain generalization, the distributions of train and test datasets could be different (for example, training on pictures of real dogs and evaluating cartoon dogs). This task requires robustness to distribution shifts and DRO is a natural framework in this context. Gulrajani & Lopez-Paz (2021) showed that the ERM framework applied over deep networks, is highly effective for this problem. Perhaps surprisingly, this remained the state-of-the-art (SOTA) algorithm for a long time. Only recently, ERM has been beaten on this challenging task (Cha et al., 2022; Addepalli et al., 2023). In this work, we show that using RGD on top of these recent techniques further boosts their performance by $0.7\%$ and gives SOTA results on this task.
Meta-Learning:
In meta-learning, we aim to learn models that generalize to new tasks with limited data. Predominant approaches in this domain use the classical ERM to optimize deep networks (Finn et al., 2017; Snell et al., 2017; Kumar et al., 2023). However, a common issue in this domain is that the learned models solve most tasks but fail catastrophically in some tasks. Consequently, this has promoted works that focus on worst-case performance (Collins et al., 2020). Recently, Li et al. (2023) used KL-DRO to tackle this problem. However, the authors applied their algorithm for solving meta-regression on a toy-dataset that is free of outliers, and haven’t showcased their algorithm on practically relevant datasets such as Omniglot, miniImageNet. In this work, we show that using RGD as an off-the-hat addition to Model-Agnostic Meta-Learning (MAML) (Finn et al., 2017) can significantly improve the worst-case accuracy of these models on meta-classification benchmarks such as Omniglot, miniImageNet by up to $3\%$.
@@ 1.2 Contributions
  • State-of-the-Art (SOTA) Performance Enhancement. Extensive experiments demonstrate that RGD delivers significant performance gains across diverse learning tasks. In tabular classification, RGD boosts the accuracy of MET (Majmundar et al., 2022) by +1.44%. For out-of-domain generalization, RGD outperforms FRR (Addepalli et al., 2023) on DomainBed by +0.7%. Additionally, RGD improves the performance of BERT on GLUE benchmarks by +1.94% and ViT on ImageNet-1K by +1.01%. (see Section 4, Appendix D)
@@ 2 Related Work
DRO dates back to the early works of Ben-Tal et al. (2009; 2013). Since then several works have studied various statistical and optimization aspects of DRO. The seminal works of Lam (2016); Namkoong & Duchi (2017); Duchi et al. (2021b) formally showed that minimizing empirical DRO risk - under small perturbations ($\rho=O(1/n)$, where $n$ is the data size) - is equivalent to minimizing sum of empirical risk and its standard deviation. Consequently, optimizing DRO leads to a better bias-variance trade-offs and generalizing models. In this work, we rely on this property of DRO to develop our re-weighting scheme. In another seminal work, Duchi & Namkoong (2018) showed that DRO - under large perturbations ($\rho=O(1)$) - leads to models with good tail performance.
@@
The aforementioned properties of DRO has led to numerous works applying it in various learning scenarios. For instance, Duchi & Namkoong (2018); Sagawa* et al. (2020); Qi et al. (2021; 2023b); Li et al. (2021; 2023) used DRO to tackle problems of class-imbalanaced classification and fairness. In another line of work, Namkoong & Duchi (2017); Fidon et al. (2020) studied DRO for designing models that generalize better than ERM. Our work falls in this second category of works that focus on generalization.
@@
From an optimization perspective, several works have focused on designing efficient algorithms for optimizing the DRO objective. These algorithms can be classified into two broad categories: primal-dual (Namkoong & Duchi, 2016; Fidon et al., 2020; Yan et al., 2020b), compositional optimization techniques (Qi et al., 2021; 2023b; 2023a; Li et al., 2021). One of the key drawbacks of primal-dual techniques is that they update and sample from a probability distribution over the entire training data at each step (aka. dual variables). A naive implementation of this step takes $O(n)$ time, which is prohibitive for large-scale tasks. Namkoong & Duchi (2016) reduced the complexity of this step to $O(\log{n})$ using data structures such as balanced binary search trees. However, the resulting algorithms are hard to implement in practice. Another drawback of these algorithms is that they require storing a buffer of weights for the entire dataset, which is infeasible at the scale of LLMs. Even if storing the weights is feasible, the presence of data augmentations complicates the resulting algorithms. Compositional optimization algorithms overcome these drawbacks by working with an equivalent dual formulation of DRO that can be written as composition of two functions: $g(\mathbb{E}_{z}[h(z,\theta)])$. One major drawback of these techniques though is that estimating the gradient $\nabla_{\theta}g(\mathbb{E}_{z}[h(z,\theta)])$ from a mini-batch is non-trivial. This requires certain additional steps in the algorithm which add to its complexity. For instance, the algorithm of Qi et al. (2021) requires making two backward passes at each step of SGD. The convergence analysis of Li et al. (2021) required two independent mini-batches at each iteration. The algorithms of Qi et al. (2023b); Li et al. (2021) both require maintaining a moving average of the weights of the mini-batches, thus adding an additional hyper-parameter to the algorithm.
Outlier robust DRO.
@@
Recent works have shown that DRO is highly sensitive to outliers (Zhai et al., 2021; Zhu et al., 2022a). This is because DRO tends to magnify the influence of outliers by upweighting them further. This is one of the reasons for the suboptimal performance of existing DRO algorithms on real world datasets. To address this, Zhai et al. (2021) consider an adversarial model for outliers (where $\epsilon$-fraction of training data could be arbitrarily corrupted by a malicious adversary). They develop a heuristic to remove the outliers during each descent step for $\chi^{2}$-DRO and CVaR. While interesting, this is too strong of an adversary model which leads to data wastage in practice.
Min-min DRO.
@@
Min-min DRO minimizes the following objective: $\inf_{P^{\prime}:D(P^{\prime}||P_{\text{data}})\leq\rho}\mathbb{E}_{P^{\prime}}[\ell(z,\theta)]$. Contrast this with DRO which minimizes: $\sup_{P^{\prime}:D(P^{\prime}||P_{\text{data}})\leq\rho}\mathbb{E}_{P^{\prime}}[\ell(z,\theta)]$. Unlike DRO which is primarily studied for generalization and fairness properties, min-min DRO is studied for training in the presence of outliers (Li et al., 2021; Kumar & Amid, 2021; Majidi et al., 2021). Instead of upweighting high loss points, min-min DRO downweights them.
Other applications of DRO.
@@
Sagawa* et al. (2020) optimized Group DRO for fair models when the group information is known. Sinha et al. (2018) studied DRO with Wasserstein divergence for learning models that are robust to adversarial perturbations. DRO also appears in many classical statistical problems. For example, many boosting algorithms (including AdaBoost) can be viewed as performing DRO with KL-divergence-based uncertainty sets (Arora et al., 2012; Friedman, 2001). Faury et al. (2020) relied on KL-DRO for counterfactual risk minimization. Sakhi et al. (2020) use DRO for improving offline contextual bandits algorithms.
Optimization Techniques for Improved Generalization.
Several optimization techniques, that fall outside the umbrella of DRO, have been proposed for improving the generalization of ML models. Of these, Sharpness-Aware Minimization (SAM) (Foret et al., 2021) is perhaps the most popular technique. From a theoretical perspective, SAM performs robust optimization in the weight space (that is SAM tries to learn a model that is robust to perturbations of weights). In contrast, RGD performs robust optimization in the distribution space. So, RGD and SAM are orthogonal to each other and can potentially be merged together to boost the performance. See Appendix D.3 for empirical comparison between RGD and SAM.